# Zero-Order Sharpness-Aware Minimization

### v1

@@ 3.3 Understanding ZOSA’s Gradient Estimation
Property 3.1 (Dang & Others, 2025) For the zeroth-order gradient estimator $\hat{g}=\frac{1}{m}\sum_{i=1}^{m}\hat{g}_{i}$, where $\hat{g}_{i}=\frac{L(\theta+\epsilon u_{i})-L(\theta)}{\epsilon}u_{i}$ and $u_{i}$ are i.i.d. Rademacher vectors in $\mathbb{R}^{d}$, under the assumption that $L(\theta)$ is twice continuously differentiable with $\mathcal{L}$-Lipschitz gradients, the expected mean squared error satisfies:
$\mathbb{E}[\|\hat{g}-\nabla L(\theta)\|^{2}]\leq\frac{\sigma^{2}}{m}+O(\epsilon),$ (7)
where $\sigma^{2}=(d-1)\|\nabla L(\theta)\|^{2}+O(\epsilon)$ bounds the variance of the loss differences, $\epsilon$ is the perturbation scale, and $m$ is the number of sampled directions. The proof is provided in Appendix A.2.
@@
Property 3.2 (Dang & Others, 2025) ZOSA employs an adaptive learning rate $\eta/\sigma_{t,\text{pert}}$, where $\sigma_{t,\text{pert}}$ is the standard deviation of the perturbed losses at the perturbed point. This design draws inspiration from normalized-SGD principles, adapting step sizes based on local curvature without the overhead of momentum. The variance of the loss perturbations satisfies:
$\mathbb{E}[\sigma_{t}^{2}]=\epsilon^{2}\|\nabla L(\theta_{t})\|^{2}+O(\epsilon^{3}),$ (8)
implying $\sigma_{t}\approx\epsilon\|\nabla L(\theta_{t})\|$. Thus, dividing by $\sigma_{t}$ normalizes the gradient estimate, yielding updates akin to:
$\theta_{t+1}=\theta_{t}-\frac{\eta}{\sigma_{t,\text{pert}}}g_{\text{pert}}\approx\theta_{t}-\eta\frac{\nabla L(\theta_{t}+\epsilon_{\text{sam}})}{\|\nabla L(\theta_{t}+\epsilon_{\text{sam}})\|}.$ (9)
This receives larger steps in flat regions (small $\sigma_{t}$) and smaller steps in steep regions, mirroring Adam-style adaptivity at inference-level memory cost. It establishes ZOSA’s equivalence to normalized-SAM in the zeroth-order domain, enhancing convergence speed and stability. The proof and Equivalence to SAM is provided in Appendix A.3.
@@ 4.1 Convergence Analysis of ZOSA
Assumption 4.2 (Bounded Variance). The stochastic gradient $\nabla L(\theta,\mathcal{B})$ has bounded variance:
$\mathbb{E}\|\nabla L(\theta,\mathcal{B})\|^{2}\leq\|\nabla L(\theta)\|^{2}+\mathcal{V}^{2},$ (11)
@@
Theorem 4.3 (Convergence of ZOSA). Assume the objective function $L(\theta)$ is $\mathcal{L}$-smooth and non-convex with a lower bound $L^{*}$, and the loss perturbations satisfy bounded variance ($\mathbb{E}[(l_{i}-L(\theta))^{2}]\leq\sigma^{2}$). The update rule is $\theta_{t+1}=\theta_{t}-\frac{\eta}{\sigma_{t,\text{pert}}}g_{\text{pert}}$, where $g_{\text{pert}}$ is the zeroth-order gradient estimate at the perturbed point $\theta_{t}+\epsilon_{\text{sam}}$, with $\epsilon_{\text{sam}}=\rho\frac{g_{t}}{\sigma_{t}}$ ($\rho>0$ is the sharpness radius), $m$ is the number of queries per estimate (Rademacher perturbations), and $\epsilon$ is the perturbation scale. Choose $\eta_{t}\leq\frac{m}{16d\mathcal{L}}$, $\epsilon\leq\frac{1}{\sqrt{d\mathcal{L}}}$, and $\rho\leq\frac{1}{4\mathcal{L}}$. Then, after $T$ iterations, ZOSA satisfies:
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla L(\theta_{t})\|^{2}\leq\frac{2(L(\theta_{1})-L^{*})}{\eta T}+2\rho\mathcal{L}+\sqrt{\frac{4d\mathcal{L}(L(\theta_{1})-L^{*})(\sigma^{2}+\epsilon^{2}G^{2})}{m\eta T}}+O(\epsilon^{2}d\mathcal{L}),$ (14)
Additionally, the algorithm biases towards approximately flat minima: for the output $\bar{\theta}$ (randomly selected from $\{\theta_{t}\}$), $\mathbb{E}[\operatorname{Tr}(\nabla^{2}L(\bar{\theta}))]\leq\min_{\theta^{*}\in\Theta^{*}}\operatorname{Tr}(\nabla^{2}L(\theta^{*}))+O(\rho+\epsilon\sqrt{d})$, where $\Theta^{*}$ is the set of minimizers and $\operatorname{Tr}$ denotes the trace of the Hessian (measuring flatness).
Remarks: ZOSA introduces the $\rho\mathcal{L}$ term for sharpness control and an explicit bias towards low-trace Hessians, enhancing generalization. The rate is $O(1/\sqrt{T})$ similar to standard ZO-SGD, but the variance term $\sqrt{d/mT}$ reflects query efficiency, with $\rho$ and $\epsilon$ terms arising from the SAM bias and perturbation scale. The proof is provided in Appendix B.1.
@@ 4.2.1 Problem Setup and Notation
Consider a training dataset $\mathcal{S}=\{(X_{i},y_{i})\}_{i=1}^{M}$ with $M$ i.i.d. samples drawn from a true data distribution $P(X,y)$. The empirical loss over $\mathcal{S}$ is defined as:
@@
$\mathbb{E}_{P(X,y)}[F(\theta;(X,y))]=\mathbb{E}_{P(X,y)}[l(\theta;(X,y))].$ (16)
@@
To enhance generalization, ZOSA adopts a sharpness-aware minimization strategy, seeking flat minima in the parameter space. We define a neighborhood around $\theta_{t}$ using the $L_{2}$ norm with an adaptive radius:
$\mathcal{C}(\theta_{t})=\left\{\delta\in\mathbb{R}^{d}\mid\|\delta\|_{2}\leq\rho/\sigma_{t}\right\},$ (18)
where $\rho>0$ controls the base neighborhood size, and the radius is adaptively scaled by $\sigma_{t}$ to account for local loss variability. The sharpness-aware objective for ZOSA is:
$\min_{\theta_{t}}\max_{\delta\in\mathcal{C}(\theta_{t})}J(\theta_{t}+\delta),$ (19)
where $J(\theta_{t}+\delta)=F(\theta_{t}+\delta;\mathcal{S})$. This formulation ensures robustness by minimizing the worst-case loss within the adaptive neighborhood.
Theorem 4.4 (Generalization Bound for ZOSA). Let the loss function $F(\theta;(X,y))$ be convex with respect to $\theta$. For any $\theta_{t}\in\mathbb{R}^{d}$ and $\sigma_{t}>0$, with probability at least $1-\kappa$ over the draw of the training set $\mathcal{S}$ with $M$ i.i.d. samples from $P(X,y)$:
$\mathbb{E}_{P(X,y)}[F(\theta_{t};(X,y))]\leq\max_{\|\delta\|_{2}\leq\rho/\sigma_{t}}F(\theta_{t}+\delta;\mathcal{S})$ (20)
$+\sqrt{\frac{k\log\left(1+\frac{\|\theta_{t}\|_{2}^{2}}{(\rho/\sigma_{t})^{2}\left(1+\sqrt{\frac{\log M}{k}}\right)^{2}}\right)+4\log\frac{M}{\kappa}+\tilde{O}(1)}{M-1}},$ (21)
where $k=d$ is the number of parameters, $\sigma_{t}$ is the adaptive standard deviation of the perturbed losses at iteration $t$ (as computed in ZOSA), and the neighborhood radius is adaptively scaled by $\sigma_{t}$. This bound links the generalization error of the parameter $\theta_{t}$ to the sharpness-aware objective and a complexity term dependent on the adaptive neighborhood size $\rho/\sigma_{t}$, dataset size $M$, parameter dimension $k$, and confidence parameter $\kappa$. The proof is provided in Appendix B.2.
@@ 5.1 Synthetic Functions
To assess the convergence of ZOSA in high-dimensional settings, we evaluated its performance on four widely used synthetic functions, including quadratic, cubic, Levy, and Rosenbrock functions (detailed definitions provided in Appendix C.1). These functions serve as standard benchmarks for optimization algorithms, enabling us to examine ZOSA’s handling of smooth, convex landscapes and challenging, ill-conditioned non-convex surfaces, where variance in zeroth-order estimates plays a critical role.
We conduct experiments in a high-dimensional regime with $d=10,000$, running for $T=40,000$ iterations. The smoothing parameter $\mu$ is set to $5^{-3}$, and the number of queries per iteration $K=1000$. For ZOSA, we use $\rho=1e-5$, $\epsilon=10^{-3}$, and $m=1000$ Rademacher vectors, tuned via grid search on a validation set. We compare ZOSA against established adaptive ZO baselines: ZO-signSGD, ZO-AdaMM, and ZO-RMSProp. All baselines share the same query budget and are initialized at $\theta_{0}$ drawn from $\mathcal{N}(0,I)$. Hyperparameters for baselines are tuned similarly, with learning rates in [1e4,0.1][1e-4,0.1] and momentum terms in $\{0.9,0.99\}$. Results are averaged over 3 independent runs.
@@
Figure 1: Convergence comparison among different adaptive ZO optimizers for various synthetic functions, in which $y$-axis represents the lg-scale optimality gap $F(\theta)-\min_{\theta^{\prime}}F(\theta^{\prime})$ and $x$-axis is the number of iterations $T$. Each curve denotes the mean from 3 independent runs.
Cosine similarity For each function, we compare the maximum (max) and average (avg) similarities between the initial gradient estimate $g_{t}$ at the original point and the sharpness-aware perturbed gradient estimate $g_{pert}$ at the perturbed point. The experimental results are presented in Fig. 3. The results demonstrate that ZOSA’s sharpness-aware mechanism, which perturbs parameters in a direction scaled by the adaptive standard deviation $\sigma_{t}$, produces perturbed gradients ($g_{pert}$-max and $g_{pert}$-avg) that exhibit competitive or improved alignment with the true gradients relative to the baseline estimates ($g_{t}$-max and $g_{t}$-avg), especially in non-convex settings such as Levy and Rosenbrock functions. This underscores the effectiveness of ZOSA in improving the stability and accuracy of gradient estimation through adaptive perturbation and variance-aware scaling, thereby enabling more robust optimization performance in challenging zero-order scenarios.
Figure 2: Cosine similarity between the estimated gradients and the true gradients across various benchmark optimization functions (Quadratic, Cubic, Levy, and Rosenbrock.
@@ 5.2 Zero-Order Prompt Fine-tuning
The zero-order prompt fine-tuning paradigm for large language models (LLMs) offers a resource-efficient pathway to tailor models for specialized tasks without gradient access or parameter exposure (Sun et al., 2022b), (Sun et al., 2023). Operating in a Language-Model-as-a-Service (LMaaS) framework, this approach relies exclusively on inference queries, making ZOSA an ideal candidate due to its sharpness-aware perturbations that target flat minima for superior generalization, combined with sigma-adaptive scaling that dynamically adjusts to estimation variance, ensuring stable and query-efficient optimization in noisy, high-dimensional prompt landscapes.
Datasets. Our evaluation spans six varied GLUE benchmarks (Wang et al., 2018), including sentiment analysis (SST-2 (Socher et al., 2013), Yelp polarity), topic classification (AG’s News (Zhang et al., 2015)), paraphrase detection (MRPC (Dolan & Brockett, 2005)), and natural language inference (RTE, SNLI (Bowman et al., 2015)). This assortment tests ZOSA’s robustness across task complexities and scales, from compact datasets like RTE ( 3.6k training samples) to expansive ones like SNLI ( 1.1M samples), showcasing its ability to handle diverse NLP demands effectively. The statistics of six datasets are summarized in Table 1. By following (Sun et al., 2022a), the testing accuracy is used to measure the performance of all the methods on the SST-2, AG’s News, RTE, SNLI, and Yelp P. datasets, and the F1 score is used to measure the performance on the MRPC datasets.
Methods. ZOSA is compared against a suite of zero-order optimizers: evolutionary algorithms including CMA-ES(Hansen, 2006) and MMES (He et al., 2020); and gradient estimators like BES(Gao & Sener, 2022), INGO (Lyu & Tsang, 2022), and SABO (Ye et al., 2024). This comparison highlights ZOSA’s superiority in fusing adaptive variance control with sharpness awareness, enabling it to excel in noisy, high-dimensional settings by minimizing estimation variance and delivering precise, targeted updates that baselines struggle to match.
Results. Table 1 displays the experimental outcomes across six benchmark datasets under three varying dimensions of the vector $v$. It is evident that the ZOSA approach surpasses all baseline methods in test classification accuracy or F1 scores in diverse configurations, underscoring its capability to enhance generalization. Remarkably, our method sustains strong performance even in the high-dimensional scenario. Fig. 3 presents a comparative analysis of ZOSA and SABO (Sharpness-Aware Black-Box Optimization) across six NLP benchmarks. Pink bars with stars represent ZOSA accuracy, green bars with circles represent SABO accuracy, blue triangles indicate ZOSA query times (calculated based on convergence iterations), and purple diamonds indicate SABO query times (computed with population size $N=100$ and 100 iterations). ZOSA, designed for rapid deployment, harnesses adaptive zeroth-order optimization techniques to achieve commendable accuracy with significantly fewer queries in certain tasks. Its query times, calculated based on actual convergence steps, exhibit variability, reflecting its efficiency in scenarios where early convergence is feasible. This suggests that ZOSA’s streamlined design could complement sharpness-aware strategies, offering a practical alternative for real-world ZO applications where query budgets are limited. Additional results for lower dimensions ($d=200$) and moderate dimensions ($d=500$), along with detailed experimental setups, are provided in Appendix C.2.

### v2

@@ 3.3 Understanding ZOSA’s Gradient Estimation
Property 3.1 (Batched One-Sided Rademacher Estimator)
For the zeroth-order gradient estimator used in ZOSA,
$\hat{g}_{t}=\frac{1}{m}\sum_{i=1}^{m}\hat{g}_{t,i},\qquad\hat{g}_{t,i}=\frac{L(\theta_{t}+\epsilon u_{i};\mathcal{B}_{t})-L(\theta_{t};\mathcal{B}_{t})}{\epsilon}u_{i},$
where $u_{1},\dots,u_{m}\overset{\text{i.i.d.}}{\sim}\text{Rademacher}^{d}$ are independent Rademacher vectors in $\mathbb{R}^{d}$, and the empirical loss $L(\cdot;\mathcal{B}_{t})$ is twice continuously differentiable with $L$-Lipschitz gradients. The second-order Hessian term vanishes exactly due to the odd symmetry of Rademacher vectors, yielding a bias of $O(\epsilon^{2})$. The variance calculation gives
$\mathbb{E}\!\left[\|\hat{g}_{t}-\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}\right]\leq\frac{(d-1)\|\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}+O(\epsilon d^{2})}{m}+O(\epsilon^{4})=O\!\left(\frac{d}{m}+\epsilon^{2}\right).$
The proof is provided in Appendix A.2.
@@
Property 3.2 (Adaptive Scaling via Loss Variance) ZOSA employs an adaptive learning rate $\eta/\sigma_{t,\text{pert}}$, where $\sigma_{t,\text{pert}}$ is the standard deviation of the perturbed losses at the perturbed point. This design draws inspiration from normalized-SGD principles, adapting step sizes based on local curvature without the overhead of momentum. The variance of the loss perturbations satisfies:
$\mathbb{E}[\sigma_{t}^{2}]=\epsilon^{2}\|\nabla L(\theta_{t};\mathcal{B}_{t})\|^{2}+O(\epsilon^{3}d),$ (9)
where $\sigma_{t}\approx\epsilon\|\nabla L(\theta_{t};\mathcal{B}_{t})\|$. Thus, dividing by $\sigma_{t}$ normalizes the gradient estimate, yielding updates akin to:
$\theta_{t+1}=\theta_{t}-\frac{\eta}{\sigma_{t,\text{pert}}}\hat{g}_{\text{pert}}\approx\theta_{t}-\eta\frac{\nabla L(\theta_{t}+\epsilon_{\text{sam}})}{\|\nabla L(\theta_{t}+\epsilon_{\text{sam}})\|}.$ (10)
This receives larger steps in flat regions (small $\sigma_{t}$) and smaller steps in steep regions, mirroring Adam-style adaptivity at inference-level memory cost. It establishes ZOSA’s approximate equivalence to normalized-SAM in the zeroth-order domain with effective sharpness radius $\rho/\epsilon$, enhancing convergence speed and stability. Detailed proofs and the approximate equivalence of SAM are provided in Appendix A.3 and Appendix A.5.
Property 3.3 (Concentration of $\sigma_{t}$) Assume the loss $L$ is $L$-smooth and $G$-Lipschitz continuous (i.e., $\|\nabla L(\theta)\|\leq G$ for all $\theta$). The perturbations $u_{i}$ are i.i.d. Rademacher vectors satisfying $\|u_{i}\|^{2}=d$, and the loss differences $\Delta l_{i}:=l_{i}-l_{0}$ are sub-Gaussian with variance proxy $V^{2}$. For batch size $m\geq\mathcal{O}(\log(1/\delta))$ (with tighter concentration in high $d$ due to CLT), with probability at least $1-\delta$,
$|\sigma_{t}-\epsilon\|\nabla L(\theta_{t})\||\leq\mathcal{O}\left(\epsilon\sqrt{d}/\sqrt{m}+\epsilon^{2}\sqrt{d}+\sqrt{\frac{V^{2}\log(1/\delta)}{m}}\right).$ (11)
This implies
$\mathrm{Var}[\sigma_{t}]\leq\mathcal{O}\left(\frac{\epsilon^{2}G^{2}+V^{2}}{m}+\frac{\epsilon^{2}\log m}{m}\right),$ (12)
ensuring that the relative variance $\mathrm{Var}[\sigma_{t}]/\mathbb{E}[\sigma_{t}]^{2}$ is small. The detailed proof is provided in Appendix A.4.
@@ 4.1 Convergence Analysis of ZOSA
Assumption 4.2 (Bounded Variance).(dang2025fzoo) The stochastic gradient $\nabla L(\theta;\mathcal{B})$ has bounded variance:
$\mathbb{E}\|\nabla L(\theta;\mathcal{B})\|^{2}\leq\|\nabla L(\theta)\|^{2}+\mathcal{V}^{2},$ (14)
@@
Theorem 4.3 (Convergence of ZOSA). Assume the objective function $L(\theta)$ is $\mathcal{L}$-smooth and non-convex with a lower bound $L^{*}$, the loss perturbations satisfy bounded variance ($\mathbb{E}[(l_{i}-L(\theta))^{2}]\leq\sigma^{2}$), and the gradients are bounded by $G$. The update rule is $\theta_{t+1}=\theta_{t}-\frac{\eta}{\sigma_{t,\text{pert}}}\hat{g}_{\text{pert}}$, where $\hat{g}_{\text{pert}}$ is the zeroth-order gradient estimate at the perturbed point $\theta_{t}+\epsilon_{\text{sam}}$, with $\epsilon_{\text{sam}}=\rho\frac{\hat{g}_{t}}{\sigma_{t}}$ ($\rho>0$ is the sharpness radius), $m$ is the number of queries per estimate (Rademacher perturbations), and $\epsilon$ is the perturbation scale. Choose $\eta_{t}\leq\frac{m}{16d\mathcal{L}},\epsilon\leq\frac{1}{\sqrt{d\mathcal{L}}},and\rho\leq\frac{1}{4\mathcal{L}}$. Then, after $T$ iterations, ZOSA satisfies:
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla L(\theta_{t})\|^{2}\leq\frac{2(L(\theta_{1})-L^{*})}{\eta T}+2\left(\frac{\rho}{\epsilon}\right)\mathcal{L}+\sqrt{\frac{4d\mathcal{L}(L(\theta_{1})-L^{*})(\sigma^{2}+\epsilon^{2}G^{2})}{m\eta T}}+O(\epsilon^{2}d\mathcal{L}),$ (17)
Additionally, the algorithm biases towards approximately flat minima: for the output $\bar{\theta}$ (randomly selected from $\{\theta_{t}\}$), $\mathbb{E}[\operatorname{Tr}(\nabla^{2}L(\bar{\theta}))]\leq\min_{\theta^{*}\in\Theta^{*}}\operatorname{Tr}(\nabla^{2}L(\theta^{*}))+O(\rho/\epsilon+\epsilon\sqrt{d})$, where $\Theta^{*}$ is the set of minimizers and $\operatorname{Tr}$ denotes the trace of the Hessian (measuring flatness).
Remarks: ZOSA introduces the $\frac{\rho\mathcal{L}}{\epsilon}$ term for sharpness control and an explicit bias towards low-trace Hessians, enhancing generalization. The rate is $O(1/\sqrt{T})$ similar to standard ZO-SGD, but the variance term $\sqrt{d/mT}$ reflects query efficiency, with $\rho$ and $\epsilon$ terms arising from the SAM bias and perturbation scale. The proof is provided in Appendix B.1.
@@ 4.2.1 Problem Setup and Notation
Consider a training dataset $\mathcal{S}=\{(X_{i},y_{i})\}_{i=1}^{M}$ with $M$ i.i.d. samples drawn from a true data distribution $\mathcal{P}(X,y)$. The empirical loss over $\mathcal{S}$ is defined as:
@@
$F(\theta)\triangleq\mathbb{E}_{(X,y)\sim\mathcal{P}(X,y)}[l(\theta;X,y)].$ (19)
@@
To enhance generalization, ZOSA incorporates a sharpness-aware minimization (SAM) strategy that seeks flat minima in the loss landscape. Following the standard SAM approximation (foret2021), we perform a single ascent step along the direction of the zeroth-order gradient estimate $\hat{g}_{t}$ to approximate the inner maximization:
$\delta_{t}=\rho\cdot\frac{\hat{g}_{t}}{\sigma_{t}}.$ (21)
[…]
This mechanism provides an efficient yet theoretically grounded approximation to the idealized sharpness-aware objective $\min_{\theta}\max_{\|\delta\|_{2}\leq\rho/\epsilon}F(\theta+\delta;\mathcal{S})$, achieving the correct effective sharpness radius without requiring gradient norm computation.
Theorem 4.4 (Non-convex Generalization Bound for ZOSA). Assume the per-sample loss $l(\cdot;\xi)$ is $L$-smooth and bounded in $[0,1]$ almost surely, and the stochastic gradient noise is $\sigma$-sub-Gaussian. Let $\theta_{T}$ be the output of ZOSA after $T$ iterations. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{M}$, we have
$\!\!F(\theta_{T})$ $\leq F(\theta_{T};\mathcal{S})+\rho\cdot\operatorname{Sharp}_{\rho/\epsilon}(\theta_{T})$
$\quad+\tilde{O}\!\left(\sqrt{\frac{D_{\mathrm{KL}}(\mathcal{N}(\theta_{T},\lambda^{-1}I)\|\mathcal{N}(0,\lambda^{-1}I))+\log(M/\delta)}{M}}+\sqrt{\frac{d\epsilon^{2}}{m}}\right)\!,$
where $\operatorname{Sharp}_{r}(\theta)\triangleq\max_{\|\delta\|_{2}\leq r}F(\theta+\delta;\mathcal{S})-F(\theta;\mathcal{S})$ is the local sharpness with radius $r$.
Remark. Although classical PAC-Bayesian bounds are inevitably loose in non-convex settings, the sharpness term $\rho\cdot\operatorname{Sharp}_{\rho/\epsilon}(\theta_{T})$ is approximately minimized by ZOSA due to its normalized SAM perturbation $\delta_{t}=\rho\cdot\hat{g}_{t}/\sigma_{t}\approx(\rho/\epsilon)\cdot\nabla F(\theta_{t};\mathcal{S})/\|\nabla F(\theta_{t};\mathcal{S})\|_{2}$. This explains ZOSA’s consistent generalization improvement on GLUE tasks. The proof is provided in Appendix B.2.
@@ 5.1 Synthetic Functions
To assess the convergence of ZOSA in high-dimensional settings, we evaluated its performance on four widely used synthetic functions, including quadratic, cubic, Levy, and Rosenbrock functions. Their specific definitions are given as follows:
$F(\bm{\theta})=\frac{1}{2}\sum_{i=1}^{d}\theta_{i}^{2},$ (22)
$F(\bm{\theta})=\sum_{i=1}^{d}|\theta_{i}|^{3}+\frac{\theta_{i}^{2}}{2},$ (23)
These functions serve as standard benchmarks for optimization algorithms, enabling us to examine ZOSA’s handling of smooth, convex landscapes and challenging, ill-conditioned non-convex surfaces, where variance in zeroth-order estimates plays a critical role.
We conduct experiments in a high-dimensional regime with d=10,000d=10{,}000, running for T=40, [^0] T=40{,}000 iterations. For the Gaussian-smoothing-based baselines (ZO-signSGD, ZO-AdaMM, and ZO-RMSProp), the smoothing parameter is set to $\mu=5\times 10^{-3}$, and the number of queries per iteration is $q=1000$. For ZOSA, we use $\rho=10^{-5}$, $\epsilon=10^{-3}$, and m=1,000m=1{,}000 batched Rademacher vectors. All methods share the same total query budget and are initialized with $\theta_{0}\sim\mathcal{N}(0,I_{d})$. Hyperparameters are tuned via grid search on a held-out validation set, with learning rates searched in $10^{-4},0.1]$ and momentum terms in $\{0.9,0.99\}$. Results are averaged over 3 independent runs with different random seeds.
@@
Figure 1: Convergence comparison among different adaptive ZO optimizers for various synthetic functions, in which $y$-axis represents the lg-scale optimality gap $F(\theta)-\text{min}_{\theta^{\prime}}F(\theta^{\prime})$ and $x$-axis is the number of iterations $T$ . Each curve denotes the mean from 3 independent runs.
Cosine similarity For each function, we compare the maximum (max) and average (avg) similarities between the initial gradient estimate $g_{t}$ at the original point and the sharpness-aware perturbed gradient estimate $g_{\text{pert}}$ at the perturbed point. The experimental results are presented in Fig. 3. The results demonstrate that ZOSA’s sharpness-aware mechanism, which perturbs parameters in a direction scaled by the adaptive standard deviation $\sigma_{t}$, produces perturbed gradients ($g_{\text{pert}}$-max and $g_{\text{pert}}$-avg) that exhibit competitive or improved alignment with the true gradients relative to the baseline estimates ($g_{t}$-max and $g_{t}$-avg), especially in non-convex settings such as Levy and Rosenbrock functions. This underscores the effectiveness of ZOSA in improving the stability and accuracy of gradient estimation through adaptive perturbation and variance-aware scaling, thereby enabling more robust optimization performance in challenging zero-order scenarios.
Figure 2: Cosine similarity between the estimated gradients and the true gradients across various benchmark optimization functions (Quadratic, Cubic, Levy, and Rosenbrock).
@@ 5.2 Zero-Order Prompt Fine-tuning
The zero-order prompt fine-tuning paradigm for large language models (LLMs) offers a resource-efficient pathway to tailor models for specialized tasks without gradient access or parameter exposure (sun2022b), (sun2023). Operating in a Language-Model-as-a-Service (LMaaS) framework, this approach relies exclusively on inference queries, making ZOSA an ideal candidate due to its sharpness-aware perturbations that target flat minima for superior generalization, combined with sigma-adaptive scaling that dynamically adjusts to estimation variance, ensuring stable and query-efficient optimization in noisy, high-dimensional prompt landscapes.
Datasets. Our evaluation spans six varied GLUE benchmarks (wang2018glue), including sentiment analysis (SST-2 (socher2013), Yelp polarity), topic classification (AG’s News (zhang2015)), paraphrase detection (MRPC (dolan2005)), and natural language inference (RTE, SNLI (bowman2015)). This assortment tests ZOSA’s robustness across task complexities and scales, from compact datasets like RTE ( 3.6k training samples) to expansive ones like SNLI ( 1.1M samples), showcasing its ability to handle diverse NLP demands effectively. The statistics of six datasets are summarized in Table 1. By following (sun2022bbtv2), the testing accuracy is used to measure the performance of all the methods on the SST-2, AG’s News, RTE, SNLI, and Yelp P. datasets, and the F1 score is used to measure the performance on the MRPC datasets.
@@
Methods. ZOSA is compared against a suite of zero-order optimizers: evolutionary algorithms including CMA-ES(hansen2006) and MMES (he2020); and gradient estimators like BES(gao2022), INGO (lyu2022), and SABO (ye2024). This comparison highlights ZOSA’s superiority in fusing adaptive variance control with sharpness awareness, enabling it to excel in noisy, high-dimensional settings by minimizing estimation variance and delivering precise, targeted updates that baselines struggle to match.
Results. Table 1 displays the experimental outcomes across six benchmark datasets under three varying dimensions of the vector $v\in\mathbb{R}^{d}$. It is evident that the ZOSA approach surpasses all baseline methods in test classification accuracy or F1 scores in diverse configurations, underscoring its capability to enhance generalization. Remarkably, our method sustains strong performance even in the high-dimensional scenario. Fig. 3 presents a comparative analysis of ZOSA and SABO (Sharpness-Aware Black-Box Optimization) across six NLP benchmarks. Pink bars with stars represent ZOSA accuracy, green bars with circles represent SABO accuracy, blue triangles indicate ZOSA query times (calculated based on convergence iterations), and purple diamonds indicate SABO query times (uses a population size of $N=100$ and is executed for $100$ iterations, following the experimental setup reported in SABO ((ye2024))). ZOSA, designed for rapid deployment, harnesses adaptive zeroth-order optimization techniques to achieve commendable accuracy with significantly fewer queries in certain tasks. Its query times, calculated based on actual convergence steps, exhibit variability, reflecting its efficiency in scenarios where early convergence is feasible. This suggests that ZOSA’s streamlined design could complement sharpness-aware strategies, offering a practical alternative for real-world ZO applications where query budgets are limited. Additional results for lower dimensions ($d=200$) and moderate dimensions ($d=500$), along with detailed experimental setups, are provided in Appendix C.2.